1)an electron moves in an electric field with a kinetic energy of 2.5ev.the associated debroglie wave length is


[h=6.626x10^-27 erg sec,me=9.108x106-28 grams,1ev=1.602x106-12 ergs]


 


 


2)the conductivity of a standard solution of MSO4 is 8.04x10^-10 Ohm^-1 cm^-1 and its equivalent conductance is 4.02 Ohm^-1 cm^2 equiv^-1.the Ksp for MSO4 will be

Grade:

1 Answers

Deepak Patra
14 years ago

Dear SRi,

Since according to the formula de broglie wavelengh is given as:

lamda = h(planks contant)/p(momentum)   i.e lamba = h/mv      --------------  eq . 1                          

where as  : p= (m)mass of electron*v(velocity of electron).

and         : K.E(kinetic enegy) = (1/2)mv2.

 so          :  v = underoot(2*K.E/m) . now sustitue value of  'v' is eq . 1.

so the debroglie wavelength: lambda = h/underoot(2*K.E*m).

Now put all the given values in the final equation

                                                    the final answer is        7.7*10-3 m.

 

askIITians.com provides online iit jee courses and IIT JEE Test Series with IITians. Click here to get free online test series and check your status timely or you can join us as our registered user for getting best iit jee study material or iit jee test series.

Click here to find IIT JEE Packages or course details with askIITians.

for further details visit the folowing links:

http://www.askiitians.com//iit-jee-online-coaching.aspx

http://www.askiitians.com/iit-jee-test-series.aspx

http://www.askiitians.com/iit-study-material/

Askiitians provide you all the best possible information, feel free to ask.

all the best.

thanks and regards.

Akhilesh Shukla